OpenStudy (anonymous):

Teaneck High had an enrollment of 11130 in 1990 and an enrollment of 11120 in year 1997. What is the predicted enrollment of Teaneck High in 2007 if we assume a linear model?

OpenStudy (anonymous):

rate of change = (11120-11130)/(1997-1990) = -10/7 year new value = 11120 -(10/7)(10) = answer